I need the 2 part solution to this absolute value problem:
|-9x+8| = |5x+8|

Thank you

Answer by Nate(3500)

|-9x + 8| = |5x + 8|
-9x + 8 = 5x + 8
-14x = 0
x = 0
also
-9x + 8 = -5x - 8
-4x = -16
x = 4